OpenStudy (anonymous):

Multiply. 5x2(2x2 + 13x – 5)

Parth (parthkohli):

\[\rm my(dog+eats-cats) = (my\times dog) + (my \times eats) - (my \times cats)\]

OpenStudy (anonymous):

XD

Parth (parthkohli):

So like...\[\rm 5x^2 (2x^2 + 13x - 5) = (5x^2 \cdot2x^2) + (5x^2\cdot13x) - (5x^2\cdot5)\]